# Efficient Model Editing with Task Vector Bases: A Theoretical Framework and Scalable Approach

### v1

@@ Abstract
Task vectors, which are derived from the difference between pre-trained and fine-tuned model weights, enable flexible task adaptation and model merging through arithmetic operations such as addition and negation. However, existing approaches often rely on heuristics with limited theoretical support, often leading to performance gaps comparing to direct task fine tuning. Meanwhile, although it is easy to manipulate saved task vectors with arithmetic for different purposes, such compositional flexibility demands high memory usage, especially when dealing with a huge number of tasks, limiting scalability. This work addresses these issues with a theoretically grounded framework that explains task vector arithmetic and introduces the task vector bases framework. Building upon existing task arithmetic literature, our method significantly reduces the memory cost for downstream arithmetic with little effort, while achieving competitive performance and maintaining compositional advantage, providing a practical solution for large-scale task arithmetic.
@@ 1 Introduction
Figure 1: Saving one task vector for each task can lead to high memory cost and suboptimal merging performance as the number of tasks increases. Instead, we can construct the affinity graph of task vectors for clustering, identify the clusters, and then apply existing merging methods on them to create task vector bases.
@@
Furthermore, the scalability of task vector-based approaches for model editing faces several challenges. One major issue is that merging models trained on multiple tasks often results in a performance drop compared to training models independently or using multi-task joint training, especially as the number of tasks increases (Ilharco et al., 2022; Yang et al., 2024). Additionally, existing model merging methods are inefficient, often requiring substantial memory to store task vectors, which are the same size as the pretrained model. For example, merging 72 fine-tuned ViT-B/32 (Dosovitskiy, 2020) models can require over 200GB of memory (Li et al., 2024; Yang et al., 2024). There were attempts that address this problem either through model sparsification (He et al., 2024; Wang et al., 2024a), alternative optimization algorithms (Li et al., 2024), or assuming all fine-tuned weights lie in a thin Gaussian shell when training task is the same (Jang et al., 2025). However, it is unclear whether the same statement applies to task arithmetic where models are finetuned from diverse tasks. This limits the applicability of such techniques, particularly when handling large-scale models or resource-constrained environments.
In this work, we address these limitations with the following contributions. First, we provide an in-depth theoretical analysis to explain when and why task arithmetic is successful. We establish the connections between key assumptions on task vector norms, task vector similarity, local smoothness, and the generalization error bounds of different task arithmetic operations. We additionally validate our assumptions and theorems with empirical evidence from previous literature and our new setups. Second, as shown in Figure 1, we leverage the similarity between different task vectors and propose a two-step framework to learn a task space with bases: first clustering the task vectors with their similarity matrix, and then merging clustered task vectors as bases with any model merging methods. Through extensive experiments, we show how to do bases arithmetic, including addition, negation and out-of-distribution task generalization with a significant reduction of memory requirement. Our task vector bases approach enables scalable and efficient model editing while preserving strong performance across multiple tasks by fully leveraging the advantage of existing task arithmetic methods in the past literature.
@@ 2 Related Work
Task arithmetic was believed to be related to the hypothesis that fine-tuning over-parameterized models behaves like a neural tangent kernel (Jacot et al., 2018), but this hypothesis contradicts the experimental results in Ortiz-Jimenez et al. (2024) where there exists a nontrivial gap between linear and standard non-linear fine-tuning generalization performance. Therefore, Ortiz-Jimenez et al. (2024) proposed a formal definition of the task arithmetic property, which has been widely adopted by subsequent work (Xiong et al., 2024) for analysis, yet under the assumption of disjoint task support, which does not hold in the standard image classification merging benchmark. Another hypothesis is related to the linear mode connectivity (Garipov et al., 2018) phenomenon (Frankle et al., 2020; Neyshabur et al., 2020), which concerns the same pretrained model finetuned on the same task with different SGD noise due to hyperparameter difference or data shuffling. The same phenomenon was observed to be layer-wise on modern model architectures (Adilova et al., 2023), and further generalized to cross-task linearity (Zhou et al., 2024b) in the context of model merging, where modes are related to different input tasks. Zhou et al. (2024b) provides a first-order Taylor expansion analysis to prove the existence of cross-task linearity relying on the convexity assumption. Similarly, the convexity assumption also appears in more recent theoretical analyses (Tao et al., 2024) that relates task arithmetic with one-shot federated learning theories, which is not required in our framework. Additionally, Zhou et al. (2024a) assumes mean-squared error as the loss function to analytically compute the optimal vector merging weights. In summary, all existing theories rely on assumptions such as disjoint task support and loss convexity that do not hold in practical models. In contrast, we do not assume loss convexity and further relax the disjoint task support assumption. We provide both discussion and empirical evidence that support our assumptions. For the literature review of other model merging methods, please refer to Appendix B.
@@
Since it is widely believed that jointly training similar tasks together improves accuracy performance or convergence (Caruana, 1993) in the multi-task learning problem, identifying task groupings is a well-explored area. Classic methods include convex formulations of task clustering (Jacob et al., 2008), latent task basis learning (Kumar & Daume III, 2012), and gradient-based methods (Fifty et al., 2021), which are more suitable for modern deep learning. When task vectors are used for addition in multi-task settings, many techniques from the multitask grouping literature can be adapted to our context. Another interesting recent work proposed the LoRA (Hu et al., 2021) Adapter library (Ostapenko et al., 2024), which uses a similar clustering approach to enable more modular large language models, along with a router-based adapter reuse algorithm. Such advantage was also proved to be successful for LoRA merging (Zhao et al., 2024), which can be seen as a special case under our bases framework. Note that we focus on the broader context of task arithmetic to examine how creating a task vector bases, or library, impacts the performance of multitask, unlearning, and domain generalization capabilities, regardless of fine tuning strategy used for task vectors.
@@ 4.1 Assumptions
Assumption 4.2 (Task Vector Near Orthogonality).
A small $\epsilon$ in 4.2 holds when any pair of tasks are not related to each other, which happens when task vectors are restricted to be sparse (He et al., 2024). Cross-task generalization when tasks are similar to each other are better understood (Tripuraneni et al., 2020; Hu et al., 2024).
Remark 4.3.
@@ 4.2 Task Arithmetic Bounds
With assumptions in the previous section, by first-order Taylor expansion w.r.t. the fine tuned model parameter, we can get the following statements for different types of task arithmetic. We defer all proof details to Appendix A.
@@
[Task Addition for Multitask Learning]theoremadd Let task addition $\theta^{*}=\theta_{0}+\sum_{i=1}^{T}\alpha_{i}\tau_{i}$ be the model parameter used for multitask learning, then $\forall i\in[T],\mathcal{L}_{i}(\theta^{*})-\mathcal{L}_{i}(\theta_{i})\leq 2L_{i}C(1+\epsilon).$ Section 4.2 shows that as long as the task vectors reside in the fine-tuning regime and task vectors are nearly orthogonal, then a single model obtained by model merging simultaneously performs comparably well on all the tasks. This bound does not depend on the number of tasks $T$. The local smoothness constant $L_{i}$ in the generalization bound implies that a flatter minima is preferred in model merging, which also partially explains the empirical success of the Fisher weighted averaging method (Matena & Raffel, 2022) as $\mathbf{H}$ is also the Fisher information matrix when $\ell$ is the cross-entropy loss which is a log-likelihood.
@@
[Out-of-Distribution Generalization]theoremgeneralize Given a collection of source task vectors $\mathcal{S}=\{\tau_{1},\tau_{2},\dots,\tau_{T}\}$ and a target task vector with $\|\tau_{\mathrm{tar}}\|^{2}\leq C$. If $\exists i\in[T]$ such that $\langle\tau_{\mathrm{tar}},\tau_{i}\rangle\geq\beta C$ for $0<\beta\leq 1$, then there exists a merging scheme $\alpha_{i},i\in[T]$ such that for the merged model $\theta^{*}=\theta_{0}+\sum_{i=1}^{T}\alpha_{i}\tau_{i},\mathcal{L}_{\mathrm{tar}}(\theta^{*})\leq\mathcal{L}_{\mathrm{tar}}(\theta_{\mathrm{tar}})+L_{\mathrm{tar}}C(1-\beta).$ This implies when $\beta$, which roughly corresponds to the cosine similarity of the two task vectors, is large enough, the gap between $\mathcal{L}_{\mathrm{tar}}(\theta^{*})$ and $\mathcal{L}_{\mathrm{tar}}(\theta_{\mathrm{tar}})$ is small, so we can use the combination of similar task vectors to achieve similar generalization performance for tasks that are out-of-distribution (OOD) w.r.t. the source models.
@@ 5 Task Vector Bases
Under limited budget constraints, it is impractical to save all task vectors for a large number of tasks $T$. Although one could argue to only save one final merged model, we lose flexibility of task vector composition, especially when we want to only merge or unlearn a part of the knowledge from a certain task in the future. Besides, the addition performance drop as the number of tasks grows is reported in Ilharco et al. (2022) in practice. Therefore, we propose using task vector bases to reduce the number of task vectors to save while retaining task information maximally and preserving the flexibility of model composition.
@@ 5.1 Bases Creation
From Section 4.2, we want $\epsilon$ to be small, thereby creating near-orthogonal bases. As observed in the green box of Figure 3, non-related tasks are nearly orthogonal, while similar task vectors represent certain interpretable skills. Based on this intuition, we define the transformation in Definition 5.1 by grouping similar task vectors using clustering algorithms to remove redundant information across tasks.
@@
In Algorithm 1, we first create the similarity matrix of task vectors and pass this matrix into spectral clustering (Ng et al., 2001) to group the tasks. Next, we can apply any existing merging method as $\texttt{merge}_{\texttt{in}}$ to create task vector bases. When cosine similarity becomes strongly negative, we first use the absolute value of the similarity for clustering and further partition the positive and negative groups with an additional clustering step. However, the scenario where task vectors have nearly opposite directions rarely occurs in practice for natural tasks unless specific adversarial tasks are defined. Therefore, lines 5-9 are typically optional. We choose spectral clustering for memory efficiency. Unlike $k$-means (Lloyd, 1982), which requires storing the entire set of $O(Td)$ task vectors during the update, spectral clustering only requires storing an $O(T^{2})$ matrix for eigenvalue computation, which is more advantageous since $d\gg T$. Additionally, in line 3 of Algorithm 1, we can determine the number of bases by inspecting the eigenspectrum of the Laplacian matrix $L_{|M|}$, without the need for additional hyperparameter tuning. This approach reduces the $T$ task vector addition problem to saving only $m$ task vector bases, significantly reducing the memory footprint.
@@ 5.2 Bases Arithmetic
We show how to do arithmetic with task vector bases in Table 1, with only slight modifications from standard task arithmetic operations discussed in Section 4.2. There are two points that need special attention. First, we observe the usage of $\texttt{merge}_{\text{out}}$, which is another iteration of model merging. We can replace $\texttt{merge}_{\text{out}}$ easily with dataless methods such as Model Soup (Wortsman et al., 2022) or TIES (Yadav et al., 2024), while for more advanced methods, we provide an example in Section 6.4.2 to show possible modifications that fit in our framework. Second, there are multiple ways to do the similarity search in Table 1. The naive method is to fine tune $\theta_{i}$ till convergence to get $\tau_{i}$ and simply compute the cosine similarity between $\tau_{i}$ and bases vectors, but then one could argue to directly use $\tau_{i}$ for unlearning and OOD generalization. Fortunately, in Section D.2, we see without training $i$ till convergence, intermediate checkpoints even with one step of gradient update still reflects important task information. Also, as we will see in Section 6.4.1 since bases can be interpreted as some skill: for example, one bases represents digit classification and the target task is MNIST, we can use such heuristics of task similarity knowledge without any explicit computation of $\tau_{i}$. In these two scenarios, since $\tau_{i}$ is either far from the optimal or does not exist, it can be more beneficial to use stored knowledge in the bases.

### v2

@@ Abstract
Task vectors, which are derived from the difference between pre-trained and fine-tuned model weights, enable flexible task adaptation and model merging through arithmetic operations such as addition and negation. However, existing approaches often rely on heuristics with limited theoretical support, often leading to performance gaps comparing to direct task fine tuning. Meanwhile, although it is easy to manipulate saved task vectors with arithmetic for different purposes, such compositional flexibility demands high memory usage, especially when dealing with a huge number of tasks, limiting scalability. This work addresses these issues with a theoretically grounded framework that explains task vector arithmetic and introduces the task vector bases framework. Building upon existing task arithmetic literature, our method significantly reduces the memory cost for downstream arithmetic with little effort, while achieving competitive performance and maintaining compositional advantage, providing a practical solution for large-scale task arithmetic. The code is available at https://github.com/uiuctml/TaskVectorBasis.
@@ 1 Introduction
Figure 1: Saving one task vector for each task can lead to high memory cost at coefficient tuning stage, especially as the number of tasks increases. Instead, we can construct the affinity graph of task vectors for clustering, identify the clusters, and then apply existing merging methods on them to create task vector bases. If we merge models with bases, coefficient tuning only happens within a strictly smaller subset of task vectors.
Figure 2: In the continual setting, we merge an incoming new task vector $\tau_{6}$ with information from the first five tasks stored in the memory bank. Storing all task vectors allows coefficient tuning for all parameters while saving single model assumes fixed relative importance within $\tau_{1}$ to $\tau_{5}$. Maintaining a single model is memory-efficient but less flexible, often leading to lower accuracy. Storing all task vectors improves accuracy but requires $O(T)$ memory. Task vector bases strike a balance, where we only constrain the weight for similar direction task vectors, reducing memory to $O(m)$ and outperforming the single model accuracy—and with proper design, sometimes even surpassing full task vectors storage’s accuracy (See Table 5, Table 8).
@@
Furthermore, the scalability of task vector-based approaches for model editing faces several challenges. One major issue is that merging models trained on multiple tasks often results in a performance drop compared to training models independently or using multi-task joint training, especially as the number of tasks increases (Ilharco et al., 2022; Yang et al., 2024). Additionally, existing model merging methods are inefficient in multiple scenarios (Figure 1, Figure 2), often requiring substantial memory to store task vectors, which are the same size as the pretrained model. For example, merging 72 fine-tuned ViT-B/32 (Dosovitskiy, 2020) models can require over 200GB of memory (Li et al., 2024; Yang et al., 2024). There were attempts that address this problem either through model sparsification (He et al., 2024; Wang et al., 2024a), alternative optimization algorithms (Li et al., 2024), or assuming all fine-tuned weights lie in a thin Gaussian shell when fine-tuning tasks are the same (Jang et al., 2025) each model. However, it is unclear whether the last two techniques applies to task arithmetic where models are finetuned from different tasks. This limits the applicability of such techniques, particularly when handling large-scale models or resource-constrained environments.
In this work, we address these limitations with the following contributions. First, we provide an in-depth theoretical analysis to explain when and why task arithmetic is successful. We establish the connections between key assumptions on task vector norms, task vector similarity, local smoothness, and the generalization error bounds of different task arithmetic operations. We additionally validate our assumptions and theorems with empirical evidence from previous literature and our new setups. Second, inspired by theoretical analysis, as shown in Figure 1, we leverage the similarity between different task vectors and propose a two-step framework to learn a task space with bases: first clustering the task vectors with their similarity matrix, and then merging clustered task vectors as bases with any model merging methods. Through extensive experiments, we show how to do bases arithmetic, including addition, negation and out-of-distribution task generalization with a significant reduction of memory requirement. Additionally, our method can be complemented with other sparsification techniques to further enhance model merging performance while maintaining low memory usage. We further improve model merging performance with lo Our task vector bases approach enables scalable and efficient model editing while preserving strong performance across multiple tasks by fully leveraging the advantage of existing task arithmetic methods in the past literature.
@@ 2 Related Work
Task arithmetic has gained attention as a computationally efficient approach for model editing and multi-task generalization (for overview empirical model merging methods, please refer to Appendix B), but its theoretical foundations remain under active investigation. Early hypotheses drew connections between task arithmetic and the neural tangent kernel (NTK) regime (Jacot et al., 2018), suggesting that fine-tuning over-parameterized models behaves linearly around the initialization. However, this hypothesis is contradicted by the empirical findings of Ortiz-Jimenez et al. (2024), which show a nontrivial generalization gap between linearized and standard fine-tuning. In response, Ortiz-Jimenez et al. (2024) proposed a formal definition of the task arithmetic property, later adopted by follow-up works such as Xiong et al. (2024), though these analyses typically assume disjoint task supports—a condition that does not hold in standard benchmarks for image classification model merging.
Compared to previous theoretical work like Ortiz-Jimenez et al. (2024), a recent work Li et al. (2025) conducts a feature-learning analysis of task arithmetic under a specific Transformer setup—namely, a one-layer, single-head architecture and a data model where labels depend on discriminative tokens. This allows analysis beyond the NTK regime and yields novel guarantees for task addition (under irrelevant or aligned tasks), task negation (under irrelevant or contradictory tasks), and OOD generalization through appropriately weighted task vector combinations. Li et al. (2025) further analyze sparsity and low-rank approximations, showing that PEFT methods do not hurt generalization guarantees. Our paper, by contrast, makes no assumption about architecture or data model. Instead, we analyze task arithmetic in a general small-norm region (4.1) around the pre-trained model, supported empirically. While our 4.4 favors near-orthogonal task relationships (analogous to their “irrelevant” case), we relax this assumption in 4.5 to demonstrate the benefits of aligned tasks for OOD generalization. We do not emphasize contradictory tasks, as such adversarial relationships are rare in naturally occurring task distributions. Our framework also inherently connects to sparsity due to the small-norm constraint, offering a complementary explanation for the generalization behavior of sparse task updates without assuming token-level feature structure.
Another line of work connects task arithmetic to the phenomenon of linear mode connectivity (Garipov et al., 2018; Frankle et al., 2020; Neyshabur et al., 2020), where models trained on the same task with different initializations or SGD noise remain connected by low-loss linear paths. This property was shown to be layer-wise in modern architectures (Adilova et al., 2023) and has been extended to cross-task linearity in the model merging setting (Zhou et al., 2024b). Theoretical support for such phenomena often relies on convexity assumptions—for example, Zhou et al. (2024b) use a first-order Taylor expansion under convex losses, and Tao et al. (2024) connect task arithmetic to one-shot federated learning using similar assumptions. Other approaches such as Zhou et al. (2024a) rely on mean-squared error objectives to derive closed-form merging weights. In contrast, our work avoids relying on loss convexity and further relaxes the disjoint task support assumption. We provide both empirical support and theoretical discussion for our assumptions, offering broader applicability to practical model merging scenarios.
@@
Since it is widely believed that jointly training similar tasks together improves accuracy performance or convergence (Caruana, 1993) in the multi-task learning problem, identifying task groupings is a well-explored area. Classic methods include convex formulations of task clustering (Jacob et al., 2008), latent task basis learning (Kumar & Daume III, 2012), and gradient-based methods (Fifty et al., 2021), which are more suitable for modern deep learning. When task vectors are used for addition in multi-task settings, many techniques from the multitask grouping literature can be adapted to our context. Another interesting recent work proposed the LoRA (Hu et al., 2021) Adapter library (Ostapenko et al., 2024), which uses a similar clustering approach to enable more modular large language models, along with a router-based adapter reuse algorithm. Such advantage was also proved to be successful for LoRA merging (Zhao et al., 2024), which can be seen as a special case under our bases framework. Note that we focus on the broader context of task arithmetic to examine how creating a task vector bases, or library, impacts the performance of multitask, unlearning, and domain generalization capabilities, regardless of fine tuning strategy (ex., parameter-efficient vs. standard fine-tuning) used for task vectors.
@@ 4.1 Common Assumptions
Smoothness is a standard assumption in optimization theory (Garrigos & Gower, 2023) and has been used in recent work on Sharpness-Aware Minimization (Foret et al., 2020; Wen et al., 2022) to encourage flatter minima and improve generalization. Since we focus mainly on the fine-tuning regime in the analysis, we only consider smoothness in a local region.
@@ 4.2 Task Arithmetic Bounds
With assumptions in the previous section, by first-order Taylor expansion w.r.t. the fine tuned model parameter, we can get the following statements for different types of task arithmetic. We defer all proof details to Appendix A. Note that two following subsections use different conditions for task similarities.
4.2.1 Addition and Negation
Assumption 4.4 (Task Similarity Upper Bound).
A small $\epsilon$ in 4.4 holds when any pair of tasks are not related to each other, which happens when task vectors are restricted to be sparse (He et al., 2024), but natural tasks can also be similar thus leading to a large $\epsilon$.
@@ 4.2.1 Addition and Negation
[Task Addition for Multitask Learning]theoremadd Let task addition $\theta^{*}=\theta_{0}+\sum_{i=1}^{T}\alpha_{i}\tau_{i}$ be the model parameter used for multitask learning, then $\forall i\in[T],\mathcal{L}_{i}(\theta^{*})-\mathcal{L}_{i}(\theta_{i})\leq L_{i}C(1+\epsilon).$ 4.4 shows that as long as the task vectors reside in the fine-tuning regime and task vectors are nearly orthogonal, then a single model obtained by model merging simultaneously performs comparably well on all the tasks. This bound does not depend on the number of tasks $T$. The local smoothness constant $L_{i}$ in the generalization bound implies that a flatter minima is preferred in model merging, which also partially explains the empirical success of the Fisher weighted averaging method (Matena & Raffel, 2022) as $\mathbf{H}$ is also the Fisher information matrix when $\ell$ is the cross-entropy loss which is a log-likelihood.
@@
4.2.2 OOD Generalization
Though similar tasks hurt addition and negation, it is possible to achieve out-of-distribution (OOD) generalization given insights from Tripuraneni et al. (2020); Hu et al. (2024).
Assumption 4.5 (Task Similarity Lower Bound).
Given a collection of source task vectors $\mathcal{S}=\{\tau_{1},\tau_{2},\dots,\tau_{T}\}$ and a target task vector with $\|\tau_{\mathrm{tar}}\|^{2}\leq C$. There exists $i\in[T]$ such that $\langle\tau_{\mathrm{tar}},\tau_{i}\rangle\geq\beta C$ for $0<\beta\leq 1$.
@@ 4.2.2 OOD Generalization
[Out-of-Distribution Generalization]theoremgeneralize There exists a merging scheme $\alpha_{i},i\in[T]$ such that for the merged model $\theta^{*}=\theta_{0}+\sum_{i=1}^{T}\alpha_{i}\tau_{i},\mathcal{L}_{\mathrm{tar}}(\theta^{*})\leq\mathcal{L}_{\mathrm{tar}}(\theta_{\mathrm{tar}})+L_{\mathrm{tar}}C(1-\beta).$ This implies when $\beta$, which roughly corresponds to the cosine similarity of the two task vectors, is large enough, the gap between $\mathcal{L}_{\mathrm{tar}}(\theta^{*})$ and $\mathcal{L}_{\mathrm{tar}}(\theta_{\mathrm{tar}})$ is small, so we can use the combination of similar task vectors to achieve similar generalization performance for tasks that are OOD w.r.t. the source models.
@@ 5 Task Vector Bases
Under limited budget constraints, it is impractical to save all task vectors for a large number of tasks $T$ for full coefficient tuning. Although one could argue to only save one final merged model, we lose flexibility of task vector composition, especially when we want to merge additional task vectors or unlearn a part of the knowledge from a certain task in the future. Therefore, we propose using task vector bases to reduce the number of task vectors to save while retaining task information maximally and preserving the flexibility of model composition.
@@ 5.1 Bases Creation
Inspired by 4.4, we aim to make $\epsilon$ small, thereby constructing near-orthogonal bases. As observed in the green box of Figure 4, unrelated tasks tend to be nearly orthogonal, while similar task vectors capture shared, interpretable skills. Based on this intuition, we define the transformation in Definition 5.1 by clustering similar task vectors to eliminate redundancy across tasks. The resulting combined similar task vectors can be roughly interpreted as “centroids” of these clusters. Since clustering encourages intra-group similarity and inter-group separation, these centroids tend to be dissimilar and thus yielding a lower $\epsilon$.
@@
In Algorithm 1, we first create the similarity matrix of task vectors and pass this matrix into spectral clustering (Ng et al., 2001) to group the tasks. Next, we can apply any existing merging method as $\texttt{merge}_{\texttt{in}}$ to create task vector bases. When cosine similarity becomes strongly negative, we first use the absolute value of the similarity for clustering and further partition the positive and negative groups with an additional clustering step. However, the scenario where task vectors have nearly opposite directions rarely occurs in practice for natural tasks unless specific adversarial tasks are defined. Therefore, lines 5-9 are typically optional. We choose spectral clustering for memory efficiency. Unlike $k$-means (Lloyd, 1982), which requires storing the entire set of $O(Td)$ task vectors during the update, spectral clustering only requires storing an $O(T^{2})$ matrix for eigenvalue computation, which is more advantageous since $d\gg T$. Additionally, in line 3 of Algorithm 1, we can determine the number of bases by inspecting the eigenspectrum of the unnormalized Laplacian matrix $L_{|M|}$, without the need for additional hyperparameter tuning. This approach reduces the $T$ task vector addition problem to saving only $m$ task vector bases, significantly reducing the memory footprint.
@@ 5.2 Bases Arithmetic
We show how to do arithmetic with task vector bases in Table 1, with only slight modifications from standard task arithmetic operations discussed in Section 4.2. There are two points that need special attention. First, we observe the usage of $\texttt{merge}_{\text{out}}$, which is another iteration of model merging. We can replace $\texttt{merge}_{\text{out}}$ easily with dataless methods such as Model Soup (Wortsman et al., 2022) or TIES (Yadav et al., 2024), while for more advanced methods, we provide an example in Section 6.4.2 to show possible modifications that fit in our framework. Second, there are multiple ways to do the similarity search in Table 1. The naive method is to fine tune $\theta_{i}$ till convergence to get $\tau_{i}$ and simply compute the cosine similarity between $\tau_{i}$ and bases vectors, but then one could argue to directly use $\tau_{i}$ for unlearning and OOD generalization. Fortunately, in Section C.2, we see without training $i$ till convergence, intermediate checkpoints even with one step of gradient update still reflects important task information. Also, as we will see in Section 6.4.1 since bases can be interpreted as some skill: for example, one bases represents digit classification and the target task is MNIST, we can use such heuristics of task similarity knowledge without any explicit computation of $\tau_{i}$. In these two scenarios, since $\tau_{i}$ is either far from the optimal or does not exist, it can be more beneficial to use stored knowledge in the bases.
Memory Requirement of Bases Arithmetic
Here we clarify the memory requirement of bases methods in two scenarios. First, in the offline setup where only one model merging operation is conducted as in Figure 1, most arithmetic methods require tuning task-specific coefficients $\alpha$ to optimize performance. This coefficient tuning phase necessitates keeping all $T$ task vectors in memory. In contrast, when using our task vector bases, if we configure both the inner and outer merging methods in Algorithm 1 as a coefficient tuning merging method, then here tuning only requires storing $\max(\max_{i}\left|\mathcal{C}_{i}\right|,m)$ vectors. By definition, this number is strictly smaller than $O(T)$, resulting in memory savings within a single merging step. Second, in the online setup like Figure 2, when a new task vector is added to an existing collection, task vector coefficients must be re-tuned for better merging performance. In this context, our approach enables storing a small set of basis vectors that can reconstruct or approximate all seen task vectors, only requiring $O(m)$ storage space of memory bank.